OpenStudy (anonymous):

is there an easier way to turn a trig function into an algebraic expression?

OpenStudy (anonymous):

Use a triangle to write \[\sin (2\sin^{-1} x)\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

no real easy way. no

OpenStudy (anonymous):

I never seem to get it..

OpenStudy (anonymous):

sorry typo there

OpenStudy (anonymous):

\[\sin(2\sin^{-1}(x))=2\sin(\sin^{-1}(x))\cos(\sin^{-1}(x))\]

OpenStudy (anonymous):

the first part is sort of obvious, \[\sin(\sin^{-1}(x))=x\]

OpenStudy (anonymous):

yes i understand that, but after that step is confusing

OpenStudy (anonymous):

they you need to know that \[\cos(\sin^{-1}(x))=\sqrt{1-x^2}\]

OpenStudy (anonymous):

which is true by any number of facts a triangle works nicely

OpenStudy (anonymous):



OpenStudy (anonymous):

or by using \(\sin^2(\theta)+\cos^2(\theta)=1\implies \cos(\theta)=\sqrt{1-\sin^2(\theta)}\implies \cos(\sin^{-1}(x))=\sqrt{1-x^2}\)